OpenStudy (taylormarie213):

John has 51 coins. Dimes and quarters worth $9.60. How many of each coin does John have?

OpenStudy (anonymous):

divide .25 (quatrer) into $9.60 dose that get yo started??

OpenStudy (anonymous):

x are dimes and y are quarters when you add all values you get 10x+ 25y= 9.60 when you add your quarters and dimes you get 51 coins x+y=51

OpenStudy (anonymous):

You now have a system of equations. You can either use substitution or addition of subtraction of the equations.

OpenStudy (anonymous):

use substitution method

OpenStudy (taylormarie213):

ah i'm braindead. Can you show me how to set up the substitution or addition of subtraction equations? please?

OpenStudy (anonymous):

i'm sorry i have to go ill try to help you later if ur answer didn't get answered but i think someone else might help you sorry

OpenStudy (anonymous):

x+y=51 isolate for x x=51-y Now replace everything on that x with 51-y 10x+ 25y= 9.60 turns into 10(51-y)+25y=9.60 Solve for y now. Once you find y use that value and plug it back in into any equation then solver for x

OpenStudy (taylormarie213):



OpenStudy (anonymous):

sorry its suppose to be .10*(51-y)+.25*y=9.60 you should get y=30

OpenStudy (taylormarie213):

thank you my friend helped me :]